# Benchmarking Legal Reasoning on 340 Law Exams

### v3

@@ Appendix A Use of AI Assistants
We use GPT-4o for coding, shortening texts and editing LaTeX more efficiently. The model is not directly used in writing, but assisting human through criticizing, grammar checks, etc.
@@ C.2 Multiple-Choice Questions
Figure 9: Model performance on open question grouped by extended features
@@ Appendix D Further Analyses on MCQs with Perturbations
We provide additional results here on MCQs with perturbations. As shown in Figure  10, across all tested settings, the models exhibit substantial variations in performance, further validating the dataset’s capability to effectively differentiate between models of varying abilities. In particular, Gemini-2.5-Pro consistently achieves the highest accuracy, maintaining a lead over the other models, while Claude-3.7-Sonnet consistently ranks second across all settings. However, all models experience a substantial drop in performance as the number of choices increases from 4 to 32. For example, the accuracy of Gemini-2.5-Pro drops from 68.6% at 4 choices to just 35.6% at 32 choices, nearly a 50% relative decrease. Similar steep declines are observed in other models: Sonnet drops from 60.9% to 33%, DeepSeek-R1 from 57.5% to 24.9%, and DeepSeek-V3 from 58.6% to 16%. The degradation is most pronounced in DeepSeek-V3, which loses over 40 percentage points in accuracy, suggesting significant difficulty with increased distractor complexity.
@@ E.4 Prompt to Answer MCQs
E.5 Sample MCQ
We provide here a sample MCQ:
{mtbenchbox}
Question:
Madeleine, a lawyer with special knowledge in insurances, is planning to open an insurance broker boutique. In order to start her business activities, she needs some advice which type of business association she should choose. Which answers are correct?
i. A company limited by shares makes the business less dependent on the person of Madeleine than a limited liability company.
ii. In a company limited by shares the company’s liability is limited to the share capital.
iii. With a company limited by shares Madeleine’s business seems more trustworthy than with a limited liability company.
iv. In order to set up a company limited by shares Madeleine needs at least CHF 100’000.
A. i, ii, iii, and iv
B. i and iii
C. none of the statements
D. iii
@@ Appendix F Inference Hyperparameters, Costs, and LLM Endpoint
For all small conventional LLMs (7 to 14B), we use vLLM [33] and one A100 GPU to do local inference. For close-sourced LLMs, we use the corresponding Official APIs. For the rest of LLMs, we use the Together AI API.
@@
Appendix G Annotation Guidelines and Additional Alt-Test Results
G.1 Annotation Protocol
All annotations were performed by three legally trained coauthors between April 30 and May 4, 2025. Each expert annotated the same 50 question–answer pairs, requiring approximately 6–8 hours in total. The annotation task was perceived as largely self-explanatory, and no extensive calibration sessions were necessary. One of the original 50 items was replaced during the process due to concerns about it not being fully self-contained.
@@ G.1 Annotation Protocol
G.2 Full Expert-Agreement Breakdown
@@
Table 10: Inter-annotator agreement among the three legal experts.
@@
[…]
G.3 Alternative Annotator Test: Additional Results
Table 11 compares three LLM judges at two tolerance levels. Deepseek-v3 fails to reach the replacement threshold ($\omega=0.00$). By contrast, both Gemini-Flash and GPT-4o comfortably exceed $\omega\geq 0.5$ under $\varepsilon=0$ and achieve perfect winning rates ($\omega=1.00$) at $\varepsilon=0.15$. Gemini-Flash attains the highest advantage probability ($\rho=0.760$), followed by GPT-4o ($\rho=0.660$), confirming their suitability as replacement judges.
Table 11: Winning rate $\omega$ and advantage probability $\rho$ for various LLM judges at two $\varepsilon$ levels. Bold $\rho$ marks the best-performing judge; underlined $\omega$ indicates $\omega\geq 0.5$.
@@ G.3 Alternative Annotator Test: Additional Results
Deepseek-v3 0.15 0.33 0.573
Gemini-2.5-Pro 0.00 1.00 0.760
Gemini-2.5-Pro 0.15 1.00 0.760
GPT-4o 0.00 0.67 0.660
GPT-4o 0.15 1.00 0.660
G.4 Alternative Annotator Test: Naive Baselines
@@
Table 12: Alternative Annotator Test results for naive constant-score baselines.
@@
Table 12 shows that constant-score baselines never achieve $\omega\geq 0.5$ at either $\varepsilon$ level. Their advantage probabilities $\rho$ peak around 0.507 for mid-range guesses (“Always 3” and “Always 4”) but remain well below the LLM judges.

### v4

@@ Appendix A Use of AI Assistants
We used AI assistants, including GPT-4o, GPT-4.5, and GPT-5, for coding, shortening texts and editing LaTeX more efficiently. The AI tools are not directly used in writing, but assisting human through criticizing, grammar checks, etc..
@@ C.2 Multiple-Choice Questions
Figure 10: Model performance on MCQ grouped by extended features
@@ Appendix D Further Analyses on MCQs with Perturbations
We provide additional results here on MCQs with perturbations. As shown in Figure  11, across all tested settings, the models exhibit substantial variations in performance, further validating the dataset’s capability to effectively differentiate between models of varying abilities. In particular, Gemini-2.5-Pro consistently achieves the highest accuracy, maintaining a lead over the other models, while Claude-3.7-Sonnet consistently ranks second across all settings. However, all models experience a substantial drop in performance as the number of choices increases from 4 to 32. For example, the accuracy of Gemini-2.5-Pro drops from 68.6% at 4 choices to just 35.6% at 32 choices, nearly a 50% relative decrease. Similar steep declines are observed in other models: Sonnet drops from 60.9% to 33%, DeepSeek-R1 from 57.5% to 24.9%, and DeepSeek-V3 from 58.6% to 16%. The degradation is most pronounced in DeepSeek-V3, which loses over 40 percentage points in accuracy, suggesting significant difficulty with increased distractor complexity.
@@ Appendix F Inference Hyperparameters, Costs, and LLM Endpoint
For all small conventional LLMs (7 to 14B), we use vLLM (Kwon et al., 2023) and one A100 GPU to do local inference. For close-sourced LLMs, we use the corresponding Official APIs. For the rest of LLMs, we use the Together AI API.
@@
Appendix G Annotation Guidelines, Judge Model Evaluations, and Full Alt-Test Results
G.1 Expert Recruitment and Evaluation Data Selection
We recruited three Swiss legal experts with PhD-level training, either in progress or completed, each of whom independently evaluated 50 examples, requiring approximately 10 hours of work per person. This substantial commitment reflects the open-ended nature of the questions, which demand both deep understanding and careful analysis of free-form texts. We selected this sample size because it aligns with the recommended leave-one-out design for the Alt-test and is a common standard in studies requiring intensive expert annotation.
The 50 questions were randomly sampled from the full pool of available questions. Although the selection was random, we ensured broad representativeness. The sample includes both German (n = 37) and English (n = 13) questions and reflects the main legal jurisdictions in similar proportions: Swiss law (n = 35), international law (n = 9), and generic law (n = 6). In addition, the sample covers 32 subdomains spanning private, public, criminal, and interdisciplinary legal areas. The distribution therefore mirrors the overall structure of the dataset.
G.2 Annotation Protocol
All annotations were carried out by three legally trained coauthors, under the supervision of a fourth coauthor who is a law professor. Each expert annotated the same 50 question–answer pairs. The annotation task was perceived as largely self-explanatory, and no extensive calibration sessions were necessary. One of the original 50 items was replaced during the process due to concerns about it not being fully self-contained.
@@ G.2 Annotation Protocol
G.3 Full Expert-Agreement Breakdown
@@
Table 10: Inter‑annotator agreement among the three legal experts.
@@
Annotator pairs show consistently strong correlation (r=0.66r=0.660.750.75), with moderate-to-substantial agreement in terms of $\kappa$. Overall, the data supports the reliability of human scores.
G.4 Experiment Design on LLM Judge and Detailed Results on Alt-Test
Previous research on LLM-as-a-Judge has identified a major drawback: judge models may exhibit self-bias or family-bias, systematically providing overly favorable evaluations of their own outputs or of outputs from models within the same family (Panickssery et al., 2024; Spiliopoulou et al., 2025). Inspired by these studies, we argue that an ideal judge model should fulfill three desiderata. First, it should achieve performance comparable to that of human experts, verified through well-established statistical methods. Second, it should be impartial, particularly when evaluating its own outputs or those from the same model family. Third, its inference should be reasonably cost-efficient, for example by using open-source models that can run at least partially locally.
We experiment using nine settings, using both closed and open models of various sizes and evaluate the results with 50 examples labeled by domain experts and summarize the full results in Table 11. We compare LLM judges at two tolerance levels. Among stand-alone models, Claude-4-Sonnet and Gemini-2.5-Pro achieve the best performance, each yielding a perfect winning rate of 1.00 in the Alt-test. In addition DeepSeek-R1 and GPT-4o pass the Alt-test with a winning rate of 0.67, indicating that they can serve as a reliable judge within our framework. However, a stand-alone judge may still exhibit self-bias or family bias, and deploying state-of-the-art commercial models entails high costs. As an alternative, we experiment with different ensemble strategies using assessments from GPT-4o, DeepSeek-V3, and Qwen3-32B—three models with relatively low costs. In particular, we find that the most effective ensembling strategy is to pool the minimum score across the three models, which significantly outperforms each model’s individual judgment quality, achieving a winning rate of 1.00 and an advantage probability of 0.713, as shown in Table 11. We include a comparison of the ensemble results and the individual assessments of the three models in Table 13.
Table 11: Winning rate $\omega$ and advantage probability $\rho$ for various LLM judges at two $\varepsilon$ levels. Bold $\rho$ and $\omega$ marks the best-performing judge models; underlined $\omega$ indicates $\omega\geq 0.5$, above the threshold to pass the test.
@@ G.4 Experiment Design on LLM Judge and Detailed Results on Alt-Test
Claude-4-Sonnet 0.00 1.00 0.780
0.15 1.00 0.780
Gemini-2.5-Pro 0.00 1.00 0.760
0.15 1.00 0.760
Ensemble 0.00 1.00 0.713
(GPT-4o & DeepSeek-V3 & Qwen3-32B) 0.15 1.00 0.713
Ensemble 0.00 0.67 0.700
(GPT-4o & DeepSeek-V3) 0.15 1.00 0.700
DeepSeek-R1 0.00 0.67 0.700
0.15 1.00 0.700
Ensemble 0.00 0.67 0.693
(GPT-4o & Qwen3-32B) 0.15 1.00 0.693
GPT-4o 0.00 0.67 0.660
0.15 1.00 0.660
Ensemble 0.00 0.00 0.607
(Deepseek-V3 & Qwen3-32B) 0.15 1.00 0.607
@@
0.15 0.33 0.573
Qwen3-32B 0.00 0.00 0.527
0.15 0.33 0.527
G.5 Alternative Annotator Test: Naive Baselines
@@
Table 12: Alternative Annotator Test results for naive constant‐score baselines.
@@
Table 12 shows that constant‐score baselines never achieve $\omega\geq 0.5$ at either $\varepsilon$ level. Their advantage probabilities $\rho$ peak around 0.507 for mid‐range guesses (“Always 3” and “Always 4”) but remain well below the LLM judges.
G.6 Full Results for Judge Model Ensembling
We present in Table 13 the full results on model performance for long-form open questions, along with bootstrapped standard errors. Qwen3-32B demonstrates self-bias, while both Qwen3-32B and GPT-4o further exhibit family bias.
Table 13: Performance of models on long-form open questions with bootstrapped standard errors. Bold numbers indicate the best scores within each group, and underlined numbers indicate the second-best scores. Wavy underlined numbers mark the best score in a row that is assigned either by the same LLM or by a model from the same family, suggesting self-bias or family bias.
Model Open Questions — Judge Score (± S.E)
Ensemble GPT-4o Deepseek-V3 Qwen3-32B
Reasoning
GPT-5 70.20 (± 0.41) 78.09 (± 0.35) 76.47 (± 0.40) 77.60 (± 0.37)
Gemini-2.5-Pro 67.40 (± 0.51) 82.23 (± 0.35) 73.95 (± 0.51) 75.74 (± 0.38)
Claude-3.7-Sonnet 62.86 (± 0.51) 77.63 (± 0.37) 69.32 (± 0.52) 71.40 (± 0.38)
GPT-5-mini 60.32 (± 0.45) 70.61 (± 0.42) 63.81 (± 0.45) 72.72 (± 0.41)
DeepSeek-R1 55.91 (± 0.51) 68.40 (± 0.49) 63.83 (± 0.50) 65.95 (± 0.43)
o3-mini 48.13 (± 0.49) 55.33 (± 0.44) 58.72 (± 0.53) 61.87 (± 0.40)
Qwen3-235B 47.25 (± 0.46) 53.40 (± 0.47) 54.24 (± 0.50) 60.11 (± 0.42)
QwQ-32B 44.36 (± 0.53) 52.86 (± 0.57) 53.27 (± 0.52) 58.20 (± 0.45)
Qwen3-32B 40.00 (± 0.43) 45.56 (± 0.46) 44.85 (± 0.43) 56.71 (± 0.43)
GPT-5-nano 27.25 (± 0.63) 31.07 (± 0.69) 29.61 (± 0.67) 34.04 (± 0.74)
Large
GPT-4.1 57.50 (± 0.51) 68.18 (± 0.40) 67.08 (± 0.53) 67.50 (± 0.42)
GPT-4o 56.93 (± 0.48) 66.24 (± 0.39) 66.09 (± 0.51) 67.41 (± 0.38)
DeepSeek-V3 52.53 (± 0.48) 59.97 (± 0.43) 62.54 (± 0.51) 64.61 (± 0.40)
Llama-4-Maverick 47.25 (± 0.46) 55.35 (± 0.39) 57.43 (± 0.51) 59.56 (± 0.38)
Llama-3.1-405B-it 43.14 (± 0.41) 48.91 (± 0.39) 53.59 (± 0.46) 55.38 (± 0.38)
Llama-3.3-70B-it 41.27 (± 0.41) 46.94 (± 0.37) 51.45 (± 0.47) 54.29 (± 0.37)
Small
GPT-4.1-mini 54.58 (± 0.43) 62.72 (± 0.42) 60.74 (± 0.46) 64.47 (± 0.40)
GPT-4.1-nano 43.68 (± 0.41) 49.46 (± 0.43) 49.27 (± 0.44) 56.74 (± 0.39)
GPT-4o-mini 42.55 (± 0.39) 48.58 (± 0.38) 52.91 (± 0.45) 55.04 (± 0.37)
Gemma-3-12B-it 41.29 (± 0.48) 50.89 (± 0.51) 48.97 (± 0.49) 55.22 (± 0.42)
Phi-4 38.54 (± 0.42) 43.45 (± 0.40) 49.80 (± 0.47) 54.03 (± 0.39)
Gemma-2-9B-it 27.41 (± 0.37) 30.75 (± 0.37) 38.96 (± 0.44) 43.23 (± 0.38)
EuroLLM-9B-it 22.95 (± 0.35) 25.68 (± 0.35) 32.30 (± 0.41) 38.17 (± 0.39)
Qwen2.5-7B-it 16.67 (± 0.29) 18.72 (± 0.29) 26.57 (± 0.37) 33.16 (± 0.36)
Ministral-8B-it 14.88 (± 0.32) 16.93 (± 0.34) 24.77 (± 0.41) 27.25 (± 0.42)
Llama-3.1-8B-it 10.00 (± 0.26) 11.40 (± 0.27) 17.62 (± 0.34) 24.79 (± 0.34)
Appendix H Quantitative Analyses on Cross-linguistic Consistency of Judge Outputs
We further conducted two analyses to quantitatively analyze the cross-lingual consistency of judge outputs.
First, we measured the Pearson correlation between four selected judge models’ scores (GPT-4o, Gemini-2.5-Pro, DeepSeek-R1, and Claude-4-Sonnet) and the average human expert scores separately for English and German questions. The experiments are conducted using the 50 examples annotated by three legal experts, the sample for the Alt-test. As summarized in Table 14, the results indicate very strong correlations for English, and slightly lower, but still strong, correlations for German. The findings are consistent across all four models. These results suggest that the models’ judgments are closely aligned with those of human experts in both languages.
Table 14: Cross-linguistic comparison of judge models
Model Pearson R (EN) P Value (EN) Pearson R (DE) P Value (DE)
GPT-4o 0.923 7.01e-06 0.761 4.67e-08
Gemini-2.5-Pro 0.908 1.79e-05 0.786 8.07e-09
DeepSeek-R1 0.937 2.30e-06 0.722 4.64e-07
Claude-4-Sonnet 0.877 8.18e-05 0.840 7.89e-11
Second, we performed a univariate linear regression to assess whether language (English vs. German) significantly explains differences between models’ and human scores. The analysis yielded non-significant effects of language, with language accounting for only a small portion of the variance, as shown in Table 15. Moreover, Gemini-2.5-Pro exhibits the lowest R², indicating highest cross-linguistic consistency. This suggests that any differences in models’ performance across languages are minimal and not statistically significant in our data. Overall, these quantitative results support the cross-linguistic robustness of the two models’ judging performance.
Table 15: Statistical tests of the language effect on the score differences
Model P Value R2
GPT-4o 0.136 0.046
Gemini-2.5-Pro 0.448 0.012
DeepSeek-R1 0.303 0.022
Claude-4-Sonnet 0.066 0.069
Appendix I Qualitative Analysis of Common Failure Modes Across Different Legal Domains
We further conducted qualitative analysis of common failure modes across different legal domains to provide insights for improving legal AI systems. Particularly, we examined models’ evaluations and found that the most frequent failure modes were (a) flawed legal reasoning and (b) references to incorrect or even non-existent laws or court decisions, and (c) weak German/multilingual proficiency. Failure modes (a) and (b) are well-documented challenges for LLMs in legal domains and appear variably across models, rather than being tied to specific questions or topics. On the other hand, failure mode (c) occurred mostly in smaller models, which often mixed languages and produced incoherent outputs.
In addition, we observed that smaller models demonstrate weaker proficiency in German and multilingual tasks compared to larger models. Some produce incoherent outputs or mix languages, resulting in nonsensical responses, an issue unique to smaller models.
We also found that GPT-4o was the most lenient judge (+1.19 points vs. human average), while Claude-4-Sonnet and Gemini-2.5-Pro were stricter (-0.39 and -0.33, respectively), and DeepSeek-R1 was moderately generous (+0.69). Higher discrepancies between model and human scores tended to occur on items where the human annotators themselves disagreed more.
Appendix J Future Plans for Data Collection
The LEXam benchmark is an important step toward evaluating long-context, process-based legal reasoning. We are preparing a second version with broader data to extend its scope, motivated by the following aspects.
LEXam includes legal questions from a wide range of subjects in both English and German. Extending the multilingual scope remains desirable, particularly by incorporating questions from jurisdictions across both common law and civil law traditions. Such expansion would enable deeper investigation into cross-system and cross-cultural differences in LLM legal reasoning, which we leave for future work.
Our data curation relies on existing courses and examination materials. Given typical timelines, legal interpretations and even applicable laws may have changed since the original questions were formulated, potentially rendering some answers outdated. Our fine-grained metadata allows easy identification of subsets affected by subsequent legal changes. Developing a mechanism to detect and update such items — similar to LiveLongBench (Wu et al., 2025) — is a promising direction.
LEXam is intended to serve not only researchers of Swiss law but also those with broader interests in legal reasoning. While this release emphasizes Swiss law, it also incorporates data from jurisdictions such as the United States, China, and the European Union. Because law varies substantially across jurisdictions, unlike universal domains such as mathematics, the creation of high-quality, multi-jurisdictional datasets is resource-intensive. To safeguard quality, we prioritized depth over breadth in the current release.
In collaboration with leading research institutions, we are expanding the benchmark to include materials from the United States, Germany, Mainland China, Hong Kong, Israel, and additional Swiss regions, with more jurisdictions planned. An open call for contributions will follow, and new data are expected in the coming months.